# An Adversarial Non-Autoregressive Model for Text Generation with Incomplete Information

### v2

@@ B.2 V-NAT and V-CMLM
Both NAT (Gu et al., 2018) and CMLM (Ghazvininejad et al., 2019) are designed for machine translation, so their original structures contains an encoder to encode input in source language. However, the input in the IIS is too limited to be processed by an encoder, so this structure can not be used in the IIS directly. Besides, the model also needs a method to sample different data. Considering the problems above, we use the idea of VAE to obtain hidden representations, so they can sample different representations and generate various sentences based on them.
@@
More specifically, a Transformer-based encoder is adopted to encode the sentences into hidden representations during training. Then, these representations are fed into the decoder to reconstruct the input sentences. During inference, representations sampled from the standard normal distribution will be fed into the decoder, and the decoder will generate sentences based on the sampled representations. The original structure of NAT and CMLM are kept as much as possible. For V-NAT the representations are fed into decoder as input (which is same with NAT). For V-CMLM, the representations are concatenated with the embeddings of input tokens (masked or unmasked words). The hyperparameters of V-NAT and V-CMLM are set as close as possible to those of ANT.
B.3 Evaluation Metrics
@@
Least Coverage Rate (LCR) (Ren and Li, 2022) is proposed to be a compliment when the FED of two models are too close to each other, since LCR is more sensitive to the change of data quality (Ren and Li, 2022). Given two sets of sentence $X_{i}^{a}\in\mathbb{X}_{a}$ and $X_{i}^{b}\in\mathbb{X}_{b}$, LCR is calculated as follows:
@@
All the token level metrics (i.e., BLEU, Self-BLEU, and Inverse BLEU) are calculated up to 5 grams following previous work (de Masson d’Autume et al., 2019; Ren and Li, 2022).

### v3

@@ A.2 Implementation of NAT and CMLM
Both NAT (Gu et al., 2018) and CMLM (Ghazvininejad et al., 2019) are designed for machine translation, so their original structures contains an encoder to encode input in source language. However, there may be no meaningful input in our experiment (e.g., unconditional generation), so this structure cannot be used in the experiment directly. We thus use the idea of VAE to obtain hidden representations, so they can be transferred to the tasks in our experiments.
@@
More specifically, a Transformer-based encoder is adopted to encode the sentences into hidden representations during training. Then, these representations are fed into the decoder to reconstruct the input sentences. During inference, representations sampled from the standard normal distribution will be fed into the decoder, and the decoder will generate sentences based on the sampled representations. For NAT, the representations are fed into decoder as input. For CMLM, the representations are concatenated with the embeddings of input tokens (masked or unmasked words). The iteration number of CMLM is set to be 10 as in previous work (Ghazvininejad et al., 2019; Huang et al., 2022b).
A.3 Evaluation Metrics
@@
Least Coverage Rate (LCR) (Ren & Li, 2023) is proposed to be a compliment when the FED of two models are too close to each other, since LCR is more sensitive to the change of data quality (Ren & Li, 2023). Given two sets of sentence $X_{i}^{a}\in\mathbb{X}_{a}$ and $X_{i}^{b}\in\mathbb{X}_{b}$, LCR is calculated as follows:
@@
All the token level metrics (i.e., BLEU, Self-BLEU, and Inverse BLEU) are calculated up to 5 grams following previous work (de Masson d’Autume et al., 2019; Ren & Li, 2023).